# Collapse-based models for gravity do not violate the entanglement-based witness of non-classicality

### v2

@@ V Discussions
In [Diosi_PhysRevA.107.062206, trillo2024diosipenrosemodelclassicalgravity, Tilloy_2016, PhysRevX.13.041040, angeli2025positivityentanglementmarkovianopen], the authors utilize a general master equation (which satisfies the completely positive map and maintains the evolution state as a classical-quantum state) to investigate the interaction between gravity and quantum matter. These equations suggest that there is a hidden quantum degree of freedom participating in the interaction process between gravity and quantum matter. While some may argue that this hidden degree of freedom is a mathematical construct rather than a physical entity, this argument is unconvincing and presents conceptual and potential causal issues.
Focusing solely on the master equation of quantum subsystems or adding correction terms only to these equations can obscure the non-classical behavior of the gravitational field or introduce a hidden degree of freedom (the third entity). Therefore, the dynamic analysis of embedding classical systems into quantum systems within a unified context is essential. Furthermore, these modified master equations should ideally be describable by a Schrödinger equation for a larger closed system, where the Hamiltonian of this larger system still comprises matter, the (extended) gravitational field, and their interaction terms. In this case, if there is a backaction from the quantum matter to the gravitational field, this implies the quantum-like nature of gravity [feng2024conservationlawsquantizationgravity].
Collapse-based models of gravity [diosi1987universal, DiosiPhysRevA.40.1165, trillo2024diosipenrosemodelclassicalgravity, Tilloy_2016, angeli2025positivityentanglementmarkovianopen] fall into the category of models that are not fundamentally semiclassical. These models introduce a hidden quantum degree of freedom in the coupling of quantum objects and the gravitational field. Our analysis and conclusions may also apply to other stochastic gravity models [PhysRevX.13.041040].
From the perspective of both theoretical predictions and experiments, true gravitationally induced entanglement can be readily distinguished from the entanglement generated by the DP model [angeli2025positivityentanglementmarkovianopen]. Consequently, the entanglement-based tests of gravity can falsify the DP model entirely [angeli2025positivityentanglementmarkovianopen, trillo2024diosipenrosemodelclassicalgravity].
@@ VI Conclusions
With the assumption of the principle of locality, gravitationally induced entanglement is the sufficient condition for the quantum nature of gravity [marletto_gravitationally_2017, bose_spin_2017, cmvvRevModPhys.97.015006], i.e., the GWT holds [cmvvRevModPhys.97.015006]. Collapse-based models of gravity, such as the DP model, do not invalidate the GWT since they somehow violate the principle of locality or introduce a hidden quantum degree of freedom to mediate the quantum information of quantum matter.
@@
Note added.– Recently, Aziz and Howl [aziz2025classical] claimed, using a quantum field–theoretic approach, that classical gravity can generate entanglement via local interactions, a claim that has prompted several refutations [marletto2025classicalgravity1, diósi2025noclassicalgravitydoes, marletto2025classicalgravity2]. Our work analyzes collapse‑based models of classical gravity and shows that, in such models, entanglement arises from a nonlocal map that cannot result from purely local gravitational interactions. This analysis supports the validity of the BMV experiment as a test of the quantum nature of gravity.
Acknowledgements.– We thank Diósi Lajos, Alan Forrester, and Matteo Carlesso for their comments on the manuscript. This research was made possible through the generous support of the Gordon and Betty Moore Foundation, the Eutopia Foundation, and of the ID 62312 grant from the John Templeton Foundation, as part of the ‘The Quantum Information Structure of Spacetime’ Project (QISS). The opinions expressed in this project/publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.

### v3

@@ V Discussions
In [10, 26, 25, 20, 1], the authors utilize a general master equation (which satisfies the completely positive map and maintains the evolution state as a classical-quantum state) to investigate the interaction between gravity and quantum matter. These equations suggest that there is a hidden quantum degree of freedom participating in the interaction process between gravity and quantum matter. While some may argue that this hidden degree of freedom is a mathematical construct rather than a physical entity, this argument is unconvincing and presents conceptual and potential causal issues.
Focusing solely on the master equation of quantum subsystems or adding correction terms only to these equations can obscure the non-classical behavior of the gravitational field or introduce a hidden degree of freedom (the third entity). Therefore, the dynamic analysis of embedding classical systems into quantum systems within a unified context is essential. Furthermore, these modified master equations should ideally be describable by a Schrödinger equation for a larger closed system, where the Hamiltonian of this larger system still comprises matter, the (extended) gravitational field, and their interaction terms. In this case, if there is a backaction from the quantum matter to the gravitational field, this implies the quantum-like nature of gravity [11].
Collapse-based models of gravity [8, 6, 26, 25, 1] fall into the category of models that are not fundamentally semiclassical. These models introduce a hidden quantum degree of freedom in the coupling of quantum objects and the gravitational field. Our analysis and conclusions may also apply to other stochastic gravity models [20].
From the perspective of both theoretical predictions and experiments, true gravitationally induced entanglement can be readily distinguished from the entanglement generated by the DP model [1]. Consequently, the entanglement-based tests of gravity can falsify the DP model and non-linear models entirely [1, 26, 13, 14].
@@ VI Conclusions
With the assumption of the principle of locality, gravitationally induced entanglement is the sufficient condition for the quantum nature of gravity [15, 4, 19], i.e., the GWT holds [19]. Collapse-based models of gravity, such as the DP model, do not invalidate the GWT since they somehow violate the principle of locality or introduce a hidden quantum degree of freedom to mediate the quantum information of quantum matter.
@@
Note added.– Recently, Aziz and Howl [2] claimed, using a quantum field–theoretic approach, that classical gravity can generate entanglement via local interactions, a claim that has prompted several refutations [18, diósi2025noclassicalgravitydoes, 16]. Our work analyzes collapse‑based models of classical gravity and shows that, in such models, entanglement arises from a nonlocal map that cannot result from purely local gravitational interactions. This analysis supports the validity of the BMV experiment as a test of the quantum nature of gravity.
Acknowledgements.– We thank Diósi Lajos, Alan Forrester, and Matteo Carlesso for their comments on the manuscript. T. Feng thanks Qiuyi Ma, Qian Chen, and Yubao Liu for their helpful discussion. This research was made possible through the generous support of the Gordon and Betty Moore Foundation, the Eutopia Foundation, and of the ID 62312 grant from the John Templeton Foundation, as part of the ‘The Quantum Information Structure of Spacetime’ Project (QISS). The opinions expressed in this project/publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.